Christensen Corporation Christensen Corporation is a wholesaler that sells a single product. Management has provided the following cost data for two levels of monthly sales volume. The company sells the product for $173.00 per unit. Sales Volume Cost of Sales Selling & Administrative Costs 4,0105.0107 $384,267 $337,173 307,567$ S317,993 1) What is the total contribution margin when 4,350 units are sold?
Answer #1
Calculation
Sales Volume 4350
Sales 752550 = 4350*173
Less: COGS 333645 = 4350*76.7
Less: Variable Costs 83433 = 4350*19.18
Contribution Margin 335472
W.N
Calculation Of Variable and Fixed Costs
Diff. in Selling And Administrative Costs = 337173-317993
19180
Variable Cost per Unit = 19180/1000
19.18